# From Landau two-fluid model to de Sitter Universe

### v9

@@ I Introduction
The factor of 2 difference between the local temperature and the temperature of the Hawking radiation from the cosmological horizon measured by the same observer has simple explanation. In case of the Hawking radiation the observer detects particles radiated from the horizon but cannot detect their partners, that are simultaneously (coherently) created on the other side of the horizon. This is similar to the doubling of the Hawking temperature of black holes suggested by ’t Hooft.Hooft2022 ; Hooft2023 ; Hooft2024 In case of the black hole horizon, the suggested partner is in the mirror image of the black hole spacetime – ”quantum clone inside black hole”.
The local temperature $T=H/\pi$ determines the local temperature in de Sitter spacetime, giving rise to the local entropy and the modified Gibbs-Duhem relation of the de Sitter state. The latter is modified due to the gravitational degrees of freedom: the gravitational coupling $K=1/16\pi G$ and the scalar Riemann curvature ${\cal R}$. These are the thermodynamically conjugate variables, which are similar to the chemical potential and particle density in condensed matter. As a result, the de Sitter can be considered as the mixture of two components: the vacuum energy (dark energy) and the artificial Zel’dovich stiff matter which plays the role of dark matter. It is the dark matter component, which is responsible for the thermodynamics of the de Sitter state with equilibrium temperature $T=H/\pi$. Together with the ordinary matter, the expanding Friedmann-Robertson-Walker Universe is represented by three components: dark energy, dark matter and conventional matter.
This is the analogue of the Landau-Khalatnikov two-fluid hydrodynamics, which we consider in Section II. This hydrodynamic theory became the source of different connections between the physics of condensed matter and relativistic physics, which includes quantum field theories, quantum gravity, and physics of the quantum vacuum. The natural consequence of the two-fluid hydrodynamics is the acoustic metric in Sec.II.1, which is produced by the ”flow of the vacuum” – the superfluid motion of the liquid. The excitations of the ”superfluid vacuum” – quasiparticles – form matter on the background of the quantum vacuum. There are different directions in the extension of the analogue gravity. Gravity emerges in the systems, which have the point nodes in the fermionic spectrum, such as the chiral superfluid 3He-A and Weyl semimetal, see Sec. II.2. Another scenario of formation of gravity comes from the elasticity tetrads, which describe deformations of crystals in Sec. II.3. In this case the quantum vacuum is considered as a superplastic crystal. Superfluid 3He-B suggests scenario, in which the gravitational tetrads emerge as the bilinear combination of fermions fields, see Sec.II.4
@@
In Section IV the two-component thermodynamics of de Sitter state is discussed. This state can be represented in terms dark energy (vacuum) and dark matter (gravitational degrees of freedom). The latter are represented by the pair of the thermodynamically conjugate variables: the gravitational coupling $K=1/16\pi G$ and the Riemann curvature ${\cal R}$. The equation of state of this dark matter is the same as Zeldovich stiff matter in Sec. IV.2 and Landau Fermi liquid in Sec. IV.3.
In Section V we consider the dynamics of de Sitter which may follow from its thermodynamics. In the suggested phenomenological approach, it is assumed that the equations are modified in such a way, that away from equilibrium the gravitational dark matter behaves as the real stiff matter. The latter experiences the cool-down due to expansion, which is partially compensated by the energy exchange with the dark energy component. The energy exchange leads to the loss of the dark energy, and as a result both dark components have the power-law decay. Both components reach at present time the correct order of magnitude. How this phenomenology is supported by microscopic theory is an open question.
@@
In Section VIII the cosmological constant problems are discussed using the experience with superfluid liquids. The main lesson from the ground states of these liquids is that the fully equilibrium quantum vacuum has exactly zero vacuum energy. This follows from the thermodynamics, which is valid both for many-body condensed matter systems and for relativistic quantum vacuum. This solves the main cosmological constant without any fine-tuning. The other cosmological problem is that it is not easy to reach the final equilibrium state. The de Sitter state has high symmetry, due to which it serves as the attractors in the dynamics of the Universe, and the final Minkowski vacuum can be obtained only by fine-tuning. However it is shown that the de Sitter vacuum is unstable. According to Section III, the de Sitter vacuum represents the heat bath for matter. This results in the thermal radiation of matter, which violates the de Sitter symmetry and thus to promotes the relaxation to the Minkowski state with zero cosmological constant.
[…]
@@ II.1 River model of quantum vacuum
In GR, the flow metric with acoustic horizon corresponds to the Painlevé-Gullstrand (PG) metric. Painleve ; Gullstrand This suggested that the gravitational field of the black hole with mass $M$ can be considered as the result of the flow of the quantum vacuum with superfluid velocity $v_{\rm s}^{2}=2GM/r$ with the event horizon at $v_{s}^{2}=c^{2}$. It is important that the PG metric is continuous across the horizon, which allows to study the thermodynamics of the vacua with horizons. Both for the real black hole and for its condensed matter analog, the Hawking radiation can be considered as semiclassical quantum tunneling across the horizon, see Ref. Volovik1999 for the analog black hole and Ref. Painleve ; Gullstrand for the real black hole.
@@ II.2 Tetrad gravity from topological Weyl point
Acoustic gravity is the part of the so-called analogue gravity, which includes also the other types of condensed matter systems, such as elastic mediaBilby1956 ; Kroener1960 ; Dzyaloshinskii1980 ; KleinertZaanen2004 ; Vozmediano2010 ; deJuan2010 ; Zaanen2010 and topological matter with Weyl fermions. The relativistic Weyl fermions in semimetals were considered by Abrikosov and BeneslavskiiAbrikosov1971 ; Abrikosov1972 ; Abrikosov1998 (both from Landau Institute). The Weyl materials allow us to simulate the horizon of black and white holes by tilting the Weyl cone.Volovik2016c ; Wilczek2020
In the fermionic Weyl and Dirac materials, emergent gravity is formulated in terms of tetrad fields, instead of the metric gravity emerging in bosonic condensed matter systems. The tetrad gravity emerges together with all the ingredients of the relativistic quantum field theories (relativistic spin, chiral Weyl fermions, gauge fields, $\Gamma$-matrices, etc.).Volovik2003 The spin of Weyl fermions in particle physics and pseudo-spin in Weyl materials forms the hedgehog in momentum space – the Weyl point in the fermionic spectrum. It represents the Berry phase monopole, which acts as a source or a sink of the Berry curvature.Volovik1987b The stability of this hedgehog is supported by topology in momentum space. It is the the topological stability of the Weyl point, which provides the emergence of the relativistic physics at low energy.
@@ II.3 Supeprlastic vacuum and translational gauge fields
The theory of elasticity in crystals can be considered in terms of elasticity tetrads.Dzyaloshinskii1980 These tetrads represent the so-called translational gauge fields, which are expressed in terms of a system of deformed crystallographic coordinate planes, surfaces of constant phase,NissinenVolovik2018b ; NissinenVolovik2018c ; NissinenVolovik2019 $X^{a}(x)=2\pi n^{a}$:
@@
It is important that such gravitational tetrads have dimensions of inverse length and inverse time, $[E^{~{}a}_{i}]=[L]^{-1}$ and $[E^{~{}a}_{0}]=[t]^{-1}$. When these elasticity tetrads are applied to general relativity (the so called superplastic vacuum KlinkhamerVolovik2019 ), one obtains that the Ricci curvature scalar ${\cal R}$ is dimensionless. The same dimensions of tetrads, $[E^{~{}a}_{i}]=[L]^{-1}$ and $[E^{~{}a}_{0}]=[t]^{-1}$, takes place in the Diakonov theory discussed in Sec. II.4.
@@ III.2 Hydrogen atom in de Sitter environment and de Sitter temperature
The ionization rate can be obtained using a simpler method.Maxfield2022 Since this process takes place well inside the horizon, one can use the classical (i.e. non-relativistic) gravitational potential $U(r)=-mv^{2}(r)/2=-mH^{2}r^{2}/2$. Then the bound state decays by quantum tunnelling of electron from the point $r=0$ to the point $r=r_{n}$, at which the electron level $-E_{n}$ matches the de Sitter gravitational potential, $U(r_{n})=-mH^{2}r_{n}^{2}/2=-E_{n}$. The radial trajectory $p_{r}(r)$ follows now from the classical equation
@@ III.6 Double Hawking temperature measured by freely-falling detector
[…]
@@ IV.1 Modified Gibbs-Duhem relation, dark energy and dark matter
The quadratic dependence of vacuum energy on temperature in Eq.(14) is important for consideration of the thermodynamic Gibbs-Duhem relation for quantum vacuum. It leads to the reformulation of the vacuum pressure. In the conventional approach, the vacuum pressure $P_{\rm vac}$ obeys the equation of state $P=w\epsilon$ with $w=-1$, i.e. $P_{\rm vac}=-\epsilon_{\rm vac}$ (let us recall that for non-relativistic matter $w=0$ and for radiation $w=1/3$). In the de Sitter state the vacuum pressure is negative, $P_{\rm vac}=-\epsilon_{\rm vac}<0$. This pressure $P_{\rm vac}$ does not satisfy the thermodynamic Gibbs-Duhem relation, $Ts_{\rm dS}=\epsilon_{\rm vac}+P_{\rm vac}$, because the right hand side of this equation is zero. The reason for that is that in this equation we did not take into account the gravitational degrees of freedom.
It was shown, that gravity contributes to thermodynamics with the pair of the thermodynamically conjugate variables: the gravitational coupling $K=\frac{1}{16\pi G}$ and the Riemann curvature ${\cal R}$, see Refs.KlinkhamerVolovik2008c ; Volovik2022G ; Volovik2020 . The gravitational thermodynamic variables allow us to write the modified Gibbs-Duhem relation:
@@
This equation allows us to introduce a pressure that characterizes the gravitational degrees of freedom and which we call the pressure of the dark matter component:
@@
Assuming that the entropy of de Sitter comes from the gravitational degrees of freedom, i.e. $s_{\rm dS}=s_{\rm DM}$, one obtains from Eq.(20) that the energy density of the dark matter equals the energy density of the vacuum, $\epsilon_{\rm DM}=\epsilon_{\rm DE}$. That is why the dark matter satisfies the equation of state $P=w\epsilon$ with $w=1$. This corresponds to stiff matter introduced by Zel’dovich,Zeldovich1962 where the speed of sound is equal to the speed of light, $c_{s}^{2}=c^{2}dP/d\epsilon_{\rm vac}=c^{2}$.
@@ IV.2 Two dark components of de Sitter and Zel’dovich stiff matter
The similar situation takes place for the closed static Universe with positive curvature introduced by Einstein. In this case there are three contributions to the pressure: from the ordinary matter, from the vacuum energy and from the gravitational degrees of freedom (i.e. from the spatial curvature $\cal R$), see Sec. 29.4.5 in the book Volovik2003 and Ref. Volovik2024e . Since the static Universe has no environment, in the full equilibrium the total pressure is zero, $P=P_{\rm vac}+P_{\rm matter}+P_{\rm gravity}=0$. The equilibrium conditions provide the connection between matter, dark energy and curvature of the Einstein Universe.
@@ IV.3 Zel’dovich stiff matter and Landau Fermi liquid
The Fermi velocity $v_{F}$ and Fermi momentum $p_{F}$ of this cosmological analog of the Fermi liquid are on the order of the speed of light and the inverse Planck length correspondingly, $v_{F}\sim c$ and $p_{F}\sim 1/l_{\rm P}\equiv M_{\rm P}/c$, where $M_{\rm P}$ is the Planck energy.
@@ V.2 Dynamics without energy exchange
$\partial_{t}\epsilon_{\rm DM}=-3H(w+1)\epsilon_{\rm DM}=-6H\epsilon_{\rm DM}\,.$ (33)
while the equation for the vacuum energy density with $w=-1$ is time independent:
@@
Of course, for the original gravitational stiff matter one has $\partial_{t}\epsilon_{\rm DM}=0$ instead of Eq.(33). But here we consider the phenomenological model, in which the real Zeldovich matter has exactly the same energy density as the gravitational Zeldovich matter in Eq.(22).
@@ V.3 Dynamics with the energy exchange
If there is the energy exchange between the two components, it is natural to assume, that these dark components tend to approach the common equilibrium state in Eq.(24). This suggests that due to the energy exchange in the process of thermalization, the energy density of dark matter tends to approach the dark energy density, $\epsilon_{\rm DM}(T)\rightarrow\epsilon_{\rm DE}(H)$. For that the equation (33) must be modified in the following way:
@@
Both components finally approach the fully equilibrium state with zero temperature – the Minkowski vacuum. That is why this phenomenology shows the route to the solution of the cosmological constant problems including the coincidence problem. All this follows from the instability of the de Sitter state in the presence of an external object (for example, a proton).
@@ VI.2 Hawking radiation as quantum tunneling
This shows that the quantum tunneling process may serve as a source of both the Hawking temperature and the Bekenstein-Hawking entropy. This is supported by consideration of macroscopic quantum tunneling, which describes the process in which a large black hole emits smaller black holes.Volovik2022G This will be discussed in Sec. VI.3.

### v10

@@ Abstract
Keywords: quantum vacuum, thermodynamics of de Sitter Universe, quantum tunneling, black holes, Unruh effect, Planck constants
@@ I Introduction
The factor of 2 difference between the local temperature and the temperature of the Hawking radiation from the cosmological horizon measured by the same observer has simple explanation. In case of the Hawking radiation the observer detects particles radiated from the horizon but cannot detect their partners, that are simultaneously (coherently) created on the other side of the horizon. This is similar to the doubling of the Hawking temperature of black holes suggested by ’t Hooft.Hooft2022 ; Hooft2023 ; Hooft2024 In case of the black hole horizon, the suggested partner is in the mirror image of the black hole spacetime – ”quantum clone inside black hole”. The factor of 2 difference between the two temperatures is also supported by topological arguments.Kusmartsev2025 ; Volovik2025f
[…]
This is the analogue of the Landau-Khalatnikov two-fluid hydrodynamics, which we consider in Section II. This hydrodynamic theory became the source of different connections between the physics of condensed matter and relativistic physics, which includes quantum field theories, quantum gravity, and physics of the quantum vacuum. The natural consequence of the two-fluid hydrodynamics is the acoustic metric in Sec.II.1, which is produced by the ”flow of the vacuum” – the superfluid motion of the liquid. The excitations of the ”superfluid vacuum” – quasiparticles – form analogue of matter on the background of the moving ”quantum vacuum”. There are different directions in the extension of the analogue gravity. Gravity emerges in the systems, which have the point nodes in the fermionic spectrum, such as the chiral superfluid 3He-A and Weyl semimetals, see Sec. II.2. Another scenario of formation of gravity comes from the elasticity tetrads, which describe deformations of crystals in Sec. II.3. In this case the quantum vacuum is considered as a superplastic crystal. Superfluid 3He-B suggests scenario, in which the gravitational tetrads emerge as the bilinear combination of fermions fields, see Sec.II.4
@@
[…]
In Section V we consider the dynamics of de Sitter state which may follow from its thermodynamics. In the suggested phenomenological approach, it is assumed that the equations are modified in such a way, that away from equilibrium the gravitational dark matter behaves as the real stiff matter. The latter experiences the cool-down due to expansion, which is partially compensated by the energy exchange with the dark energy component. The energy exchange leads to the loss of the dark energy, and as a result both dark components have the power-law decay. Both components reach at present time the correct order of magnitude. How this phenomenology is supported by microscopic theory is an open question.
@@
In Section VIII the cosmological constant problems are discussed using the experience with superfluid liquids. The main lesson from the ground states of these liquids is that the fully equilibrium quantum vacuum has exactly zero vacuum energy. This follows from the thermodynamics, which is valid both for many-body condensed matter systems and for relativistic quantum vacuum. This solves the main cosmological constant without any fine-tuning. The other cosmological problem is that it is not easy to reach the final equilibrium state. The de Sitter state has high symmetry, due to which it serves as the attractors in the dynamics of the Universe, and the final Minkowski vacuum can be obtained only by fine-tuning. However it is shown that the de Sitter vacuum is unstable. According to Section III, the de Sitter vacuum represents the heat bath for matter. This results in the thermal radiation of matter, which violates the de Sitter symmetry and thus promotes the relaxation to the Minkowski state with zero cosmological constant.
In Section IX we discuss the dynamic origin of Planck constants $\hbar$ and $h=\hbar c$ as elements of the tetrads in Minkowski spacetime. We also discuss the space-time dimensions of the physical quantities. In particular the Planck constants $\hbar$ and /h/\!\!\!h have dimensions of time and space correspondingly, and thus they cannot be considered as the fundamental constants. The effective Planck constant $h_{\rm ac}$ emerging in the acoustic gravity in superfluid 4He has dimension of length and the size of the interatomic distance. Some dimensionless quantities can be quantized, which is considered on example of the entropies of the cosmological and black hole horizons.
@@ II.1 River model of quantum vacuum
In GR, the flow metric with acoustic horizon corresponds to the Painlevé-Gullstrand (PG) metric. Painleve ; Gullstrand This suggested that the gravitational field of the black hole with mass $M$ can be considered as the result of the flow of the quantum vacuum with superfluid velocity $v_{\rm s}^{2}=2GM/r$ with the event horizon at $v_{s}^{2}=c^{2}$. It is important that the PG metric is continuous across the horizon, which allows to study the thermodynamics of the vacua with horizons. Both for the real black hole and for its condensed matter analog, the Hawking radiation can be considered as semiclassical quantum tunneling across the horizon, see Ref. Volovik1999 for the analog black hole and Refs. ParikhWilczek2000 ; Srinivasan1999 for the real black hole.
@@ II.2 Tetrad gravity from topological Weyl point
Acoustic gravity is the part of the so-called analogue gravity, which includes also the other types of condensed matter systems, such as elastic mediaBilby1956 ; Kroener1960 ; Dzyaloshinskii1980 ; KleinertZaanen2004 ; Vozmediano2010 ; deJuan2010 ; Zaanen2010 and topological matter with Weyl fermions. The relativistic Weyl fermions in semimetals were considered by Abrikosov and BeneslavskiiAbrikosov1971 ; Abrikosov1972 ; Abrikosov1998 (both from Landau Institute). The Weyl materials allow us to simulate the horizon of black and white holes by tilting the Weyl cone.Volovik2016c ; Wilczek2020 ; Baskaran2025
In the fermionic Weyl and Dirac materials, emergent gravity is formulated in terms of tetrad fields, instead of the metric gravity emerging in bosonic condensed matter systems. The tetrad gravity emerges together with all the ingredients of the relativistic quantum field theories (relativistic spin, chiral Weyl fermions, gauge fields, $\Gamma$-matrices, etc.).Volovik2003 The spin of Weyl fermions in particle physics and pseudo-spin in Weyl materials forms the hedgehog in momentum space – the Weyl point in the fermionic spectrum. It represents the Berry phase monopole, which acts as a source or a sink of the Berry curvature.Volovik1987b The stability of this hedgehog is supported by topology in momentum space. It is the topological stability of the Weyl point, which provides the emergence of the relativistic physics at low energy.
@@ II.3 Supeprlastic vacuum and translational gauge fields
The theory of elasticity in crystals can be considered in terms of elasticity tetrads.Dzyaloshinskii1980 These tetrads represent the so-called translational gauge fields, which are expressed in terms of a system of deformed crystallographic coordinate planes, surfaces of constant phase,NissinenVolovik2018b ; NissinenVolovik2019 ; NissinenHeikkila2021 ; Nissinen2022 $X^{a}(x)=2\pi n^{a}$:
@@
[…]
@@ III.2 Hydrogen atom in de Sitter environment and de Sitter temperature
The ionization rate can be obtained using a simpler method.Maxfield2022 Since this process takes place well inside the horizon, one can use the classical (i.e. non-relativistic) gravitational potential $U(r)=-mv^{2}(r)/2=-mH^{2}r^{2}/2$. Then the bound state decays by quantum tunnelling of electron from the point $r=0$ to the point $r=r_{n}$, at which the electron level $E_{n}$ matches the de Sitter gravitational potential, $U(r_{n})=-mH^{2}r_{n}^{2}/2=-E_{n}$. The radial trajectory $p_{r}(r)$ follows now from the classical equation
@@ III.6 Double Hawking temperature measured by freely-falling detector
In the case of Hawking radiation, we have $r_{1}<r_{H}<r_{2}$, and the integral $\int dr\,\delta(r-r_{H})=1$. As a result, we obtain the conventional Hawking temperature, $T_{H}=|dv/dr|_{r=r_{H}}/2\pi$. In the case of the detector crossing the horizon, the radiation is caused by the detector itself. The detector plays the role of an external object that produces radiation in the same way as an external atom provides radiation of an electron in a de Sitter environment. In this case, $r_{1}$ is determined by the position of the detector, $r_{1}=r_{H}<r_{2}$, and the integral is $\int dr\,\delta(r-r_{H})=1/2$. As a result, the local temperature measured by the detector is twice as large, $T=|dv/dr|_{r=r_{H}}/\pi=2T_{H}$.
@@ IV.1 Modified Gibbs-Duhem relation, dark energy and dark matter
The quadratic dependence of vacuum energy on temperature in Eq.(14) is important for consideration of the thermodynamic Gibbs-Duhem relation for quantum vacuum. It leads to the reformulation of the vacuum pressure. In the conventional approach, the vacuum pressure $P_{\rm vac}$ obeys the equation of state $P=w\epsilon$ with w=1w=-1, i.e. $P_{\rm vac}=-\epsilon_{\rm vac}$ (let us recall that for non-relativistic matter $w=0$ and for radiation $w=1/3$). In the de Sitter state the vacuum pressure is negative, $P_{\rm vac}=-\epsilon_{\rm vac}<0$. This pressure $P_{\rm vac}$ does not satisfy the thermodynamic Gibbs-Duhem relation, $Ts_{\rm dS}=\epsilon_{\rm vac}+P_{\rm vac}$, because the right hand side of this equation is zero. The reason for that is that in this equation we did not take into account the gravitational degrees of freedom.
[…]
@@
This equation allows us to introduce a pressure that characterizes the gravitational degrees of freedom:
@@
We call $P_{DM}$ the pressure of the dark matter component, although it has nothing to do with real physical dark matter.
Assuming that the entropy of de Sitter comes from the gravitational degrees of freedom, i.e. $s_{\rm dS}=s_{\rm DM}$, one obtains from Eq.(20) that the energy density of this ‘dark matter’ equals the energy density of the vacuum, $\epsilon_{\rm DM}=\epsilon_{\rm DE}$. That is why our dark matter satisfies the equation of state $P=w\epsilon$ with $w=1$. This corresponds to stiff matter introduced by Zel’dovich,Zeldovich1962 where the speed of sound is equal to the speed of light, $c_{s}^{2}=c^{2}dP/d\epsilon_{\rm vac}=c^{2}$.
@@ IV.2 Two dark components of de Sitter and Zel’dovich stiff matter
[…]
@@ IV.3 Zel’dovich stiff matter and Landau Fermi liquid
The Fermi velocity $v_{F}$ and Fermi momentum $p_{F}$ of this cosmological analog of the Fermi liquid are on the order of the speed of light and the inverse Planck length correspondingly, $v_{F}\sim c$ and $p_{F}\sim 1/l_{\rm P}\equiv M_{\rm P}/c$, where $M_{\rm P}$ is the reduced Planck mass, $M_{\rm P}^{2}=1/(8\pi G)$.
@@ V.2 Dynamics without energy exchange
$\partial_{t}\epsilon_{\rm DM}=-3H(w+1)\epsilon_{\rm DM}=-6H\epsilon_{\rm DM}\,$, (33)
while the equation for the vacuum energy density with w=1w=-1 is time independent:
@@
Of course, for the original gravitational stiff matter one has $\partial_{t}\epsilon_{\rm DM}=0$ instead of Eq.(33). But here we consider the toy phenomenological model, in which the real Zeldovich matter has exactly the same energy density as the gravitational Zeldovich matter in Eq.(22).
@@ V.3 Dynamics with the energy exchange
If there is the energy exchange between the two components, it is natural to assume, that these dark components tend to approach the common equilibrium state in Eq.(24). This suggests that due to the energy exchange in the process of thermalization, the temperature of the dark matter tends to approach the temperature $T=H/\pi$ of the de Sitter thermal bath, and thus the energy density of dark matter tends to approach the dark energy density, $\epsilon_{\rm DM}(T)\rightarrow\epsilon_{\rm DE}(H)$. For that the equation (33) must be modified in the following way:
@@
Both components finally approach the fully equilibrium state with zero temperature – the Minkowski vacuum. That is why this toy phenomenological model shows the route to the solution of the cosmological constant problems including the coincidence problem (see also Section VIII.3). The key point here is the instability of the de Sitter state in the presence of an external object (for example, a proton).
@@ VI.2 Hawking radiation as quantum tunneling
This shows that the quantum tunneling process may serve as a source of both the Hawking temperature and the Bekenstein-Hawking entropy. This is supported by consideration of macroscopic quantum tunneling discussed in Sec. VI.3. It describes the process in which a large black hole emits smaller black holes.Volovik2022G
Moreover, the macroscopic quantum tunneling from the black hole to the white hole with the same mass demonstrates the curious thermodynamic properties of the white hole.Volovik2025P In particular, the entropy of the white hole is negative, $S_{\rm WH}(M)=-S_{\rm BH}(M)$. This is a consequence of a special type of event horizon that also occurs in a collapsing de Sitter universe with $H<0$. It is unclear whether negative entropy can arise in systems without an event horizon and in condensed matter analogues.